OpenStudy (anonymous):

Stats: In a large population of high school students, 20% have experienced math anxiety. You take a random sample of 10 students from this population. The standard deviation of the number of students in the sample who have experienced math anxiety is?? :)

OpenStudy (mathmale):

Breee: I recognize this problem. I recommended that you try to find the proper formula for sample standard deviation when working with proportions. What did you find through your Internet search for that topic?

OpenStudy (mathmale):

My point here, Bree, is that developing the ability to find needed info is essential, the higher you go in education. I will share my own Internet search for "proportion standard deviation." https://www.google.com/search?sourceid=chrome-psyapi2&ion=1&espv=2&es_th=1&ie=UTF-8&q=proportion%20sample%20standard%20deviation&oq=proportion%20sample%20standard%20deviation&aqs=chrome..69i57j0l3.9336j0j7

OpenStudy (anonymous):

sqrt((p(1-p)/n) is what I found but I'm not sure what to do...

OpenStudy (mathmale):

You're getting there. Please refer to the following: http://sites.stat.psu.edu/~dhunter/100/spring2005/lecture/lecture28.pdf There are 4 boxes with info in them at the top of this page. Look at the box in the 2nd column and the second row. This is the formula you need for "sample standard deviation" for a sample proportion.

OpenStudy (mathmale):

In your particular problem, you may use p=0.20 and n=10 (since you have 10 samples).

OpenStudy (mathmale):

If you can evaluate that formula correctly, you're done.

OpenStudy (mathmale):

copy down the formula from http://sites.stat.psu.edu/~dhunter/100/spring2005/lecture/lecture28.pdf onto paper and then subst. p=0.2 and n=10. share your work, if possible. Your result??

OpenStudy (zarkon):

let \(X\) be the number of students out of the 10 sampled that have math anxiety Then \(X\sim\)Bin(10,.2) the s.d. of \(X\) is given by \(\sqrt{np(1-p)}\)

OpenStudy (mathmale):

Zarkon: I know there are criteria for approximating the sample standard deviation in this manner, but have forgotten them. Are they worth bringing up in this discussion? Thank you for the correction to my statement earlier.

OpenStudy (zarkon):

you are not looking for the standard deviation of the proportion \[\frac{X}{n}\] which would be \[\sqrt{\frac{p(1-p)}{n}}\] you are looking for the s.d. of \(X\)

OpenStudy (anonymous):

Sorry! My computer crashed :o

OpenStudy (anonymous):

So it would be sqrt((.20(1-.20))/10)?

OpenStudy (zarkon):

no...use the formula \[\sqrt{np(1-p)}\]

OpenStudy (anonymous):

Oh okay one sec

OpenStudy (anonymous):

sqrt(10*.2(1-.2)) = 1.265?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Both equations gave me the same answer, what's the difference between the two??

OpenStudy (zarkon):

they don't...you must have put it in your calculator incorrectly

OpenStudy (mathmale):

Thanks again, Zarkon. Bree: Sorry for having given you the wrong formula.

OpenStudy (anonymous):

Okey dokey, thanks for the help! :)

OpenStudy (anonymous):

It's all good :)

OpenStudy (zarkon):

the previous formula gives .1265 the latter is 1.265

OpenStudy (anonymous):

Ohhh oops, my bad :)

OpenStudy (anonymous):

Thanks again! :D